# pulsars; FRBs

### 08 Apr 2024 v1

@@ 3 A spectrum of solutions
Figure 3: Intermediate solutions for $\theta_{\rm pc}=\pi/2$ and $\theta_{\rm pc}=2/1.176/0.9/0.7/0\times(r_{*}/R_{\rm LC})^{1/2}$ (top left/top right/middle left/middle right/bottom left/bottom right respectively), or equivalently $\Psi_{\rm open}=\Psi_{\rm max}$ and $\Psi_{\rm open}\equiv\Psi_{\rm S}=\Psi_{\rm max}\sin^{2}\theta_{\rm pc}=2.256/1.228/0.756/0.468/0\times\Psi_{\rm dipole\ LC}$ respectively. $\Psi$ contour levels: integer multiples of $0.1\Psi_{\rm dipole\ LC}$ with $\Psi=0$ along the $z$-axis. Red dotted lines: initial dipolar shape of the separatrix between open and closed field lines. Thin vertical line: line cylinder. Thick lines: current sheets along the separatrix and equator (and along the light cylinder in the bottom right solution).
@@
We obtain here magnetospheric solutions with the methodology of Paper I (see Appendix 1) in which the separatrix between open and closed field lines is considered to be a mathematical contact discontinuity, namely a surface of zero thickness. The tip of the closed-line region is called the Y-point. We consider in particular Y-point positions from $x_{\rm Y}\approx r_{*}/R_{\rm LC}\ll 1$ to $x_{\rm Y}\approx 1$ very close to the light cylinder. For reasons of computational convenience we chose a rather large star with $r_{*}=0.25R_{\rm LC}$ (PINN methods cannot handle well large scale differences in their domain of application). We have also tried different (smaller) values of $r_{*}$ and obtained similar results. As is described in Appendix 1, we first choose the angular opening of the polar cap $\theta_{\rm pc}$ from which a certain amount of magnetic flux $\Psi_{\rm open}\equiv\Psi_{\rm S}=\Psi_{\rm max}\sin^{2}\theta_{\rm pc}$ emanates and extends to infinity. This forms the so-called open-line region. In the solutions shown in figure 3, one particular value of $\theta_{\rm pc}$, namely $\theta_{\rm pc}=1.176(r_{*}/R_{\rm LC})^{1/2}$, was specially chosen with the following in mind: It is straightforward to calculate that, in a dipolar magnetic field configuration, the magnetic field line that crosses the light cylinder corresponds to $\Psi_{\rm dipole\ LC}=\Psi_{\rm max}r_{*}/R_{\rm LC}=B_{*}r_{*}^{3}\Omega/(2c)$ (here, $B_{*}$ is the polar value of the surface magnetic field), and $\theta_{\rm dipole\ pc}=\sin^{-1}[(r_{*}/R_{\rm LC})^{1/2}]\approx(r_{*}/R_{\rm LC})^{1/2}$. In previous high-resolution solutions of the pulsar equation (e.g. Timokhin (2006)), the magnetic field line that crosses the light cylinder corresponds to $\Psi_{\rm LC\ Timokhin}=1.23\Psi_{\rm dipole\ LC}$ and $\theta_{\rm pc\ Timokhin}\approx 1.176(r_{*}/R_{\rm LC})^{1/2}$. Therefore, this particular value of $\theta_{\rm pc}$ was chosen so that the Y-point lies very close to the light cylinder. Instead, in Paper I it was found that $R_{\rm Y}=0.88R_{\rm LC}$ which is closer to $R_{\rm dipole}(\Psi_{\rm S})=r_{*}/\sin^{2}\theta_{\rm pc}=0.81R_{\rm LC}$. This unexpected result is due to our clear treatment of the separatrix surface as a surface of zero thickness which allows the manifestation of the T-point at its tip (see Appendix 2). This is not the case in all previous MHD, FFE, and PIC simulations where the separatrix surface has an unphysically large finite thickness which at its tip leads to an exterior Y-point, and a T-point some distance inside it (see figure 3 of Contopoulos et al. (2024) for details).
@@
In figure 3 it is also shown the solution for $\theta_{\rm pc}=0.868(r_{*}/R_{\rm LC})^{1/2}$ that yields $R_{\rm Y}=0.91R_{\rm LC}$ and $\Psi_{\rm open}=0.71\Psi_{\rm dipole\ LC}$. The reader can check that this value of $\Psi_{\rm open}$ corresponds to a pulsar spindown rate equal to $\dot{E}=0.75\dot{E}_{\rm vacuum}(90^{\circ})$. Here, $\dot{E}_{\rm vacuum}(\lambda)$ is the spindown rate of a vacuum dipole rotator with inclination angle $\lambda$, and $\dot{E}_{\rm vacuum}(90^{\circ})=B_{*}^{2}r_{*}^{6}\Omega^{4}/(6c^{3})\equiv 2\Psi_{\rm dipole\ LC}^{2}\Omega^{2}/(3c)$. One may tentatively generalize our result for non-zero pulsar inclination angles according to Spitkovsky (2006) as $\dot{E}(\lambda)\approx 0.75\dot{E}_{\rm vacuum}(90^{\circ})(1+\sin^{2}\lambda)$, and since $\dot{E}_{\rm vacuum}(\lambda)=\dot{E}_{\rm vacuum}(90^{\circ})\sin^{2}\lambda$, we obtain that
@@
It is interesting that in all previous solutions of the FFE pulsar magnetosphere, the above ratio was found to be greater than 3 (e.g. Li et al. (2012)). This value is significantly larger than the ratio of spindown rates $\dot{E}_{\rm ON}/\dot{E}_{\rm OFF}$ observed in the intermittent pulsars PSR B1931+24, PSR J1832+0029 and PSR J1841-0500 for their correponding ‘ON’ and ‘OFF’ states (1.5, 1.7 and 2.5 respectively; e.g. Rea et al. (2008); Wang et al. (2020)). The inability to account for observed values lower than 3 was the reason that led to the development of resistive magnetospheric solutions (e.g. Kalapotharakos et al. (2012); Li et al. (2012)). With our new solutions it seems that there is no need for magnetospheric resistivity anymore. This result certainly merits further investigation.
We were able to apply our method to even smaller values of $\Psi_{\rm open}$ and $\theta_{\rm pc}$. These solutions are very hard to obtain with the methodology of Paper I because in these the Y-point approaches very close to the light cylinder and the convergence of the solution in the open-line region between the Y-point and the light cylinder is very difficult. In the limit $\Psi_{\rm open}=\theta_{\rm pc}=0$ one reaches the solution shown in figure 1 which was obtained independently as a solution of the pulsar equation (eq. 2) with standard methods (see Gourgouliatos and Lynden-Bell (2008)). This spectrum of solutions with smaller and smaller spindown rates is new and has never been obtained before in previous MHD, FFE, nor PIC simulations. Our results are summarized in Table 1.
@@
In figure 4 the distribution of $B_{z}$ is plotted along the equator inside the light cylinder and show its clear increase right inside the Y-point at its tip, as is needed in order to satisfy eq. (13). Although this effect was known, it has never been observed so clearly before, and figure 4 improves figure 11 of Timokhin (2006). It is now seen clearly that, with the improved methodology of Dimitropoulos et al. (2024), the minimum $B_{z}$ value of this sequence of solutions lies at around $x_{\rm Y}\sim 0.8$ which is near the position of the Y-point in the high-resolution PIC simulations of Cerutti et al. (2015); Hu and Beloborodov (2022); Hakobyan et al. (2023); Bransgrove et al. (2023). It may be that the Y-point naturally relaxes to that position (not at the light cylinder), as this corresponds to a minimum energy solution. This result may answer an important question in the recent literature on the pulsar magnetosphere. Note, however, that the numerical experiments in Kalapotharakos et al. (2023) lead to a different interpretation, namely that the Y-point can be pushed arbitrarily close to the stellar surface depending on the number of PIC particles that are either injected or reach the dissipative region beyond the Y-point. This is an interesting interpretation since it suggests that $x_{\rm Y}$ may decrease as the magnetosphere evolves with pulsar spindown. This will lead to a value of the braking index smaller than its canonical value of 3. Otherwise, if $x_{\rm Y}$ remains unchanged, all solutions obtained in Table 1 will have braking indices equal to 3.
@@
In figure 5 the variation of the open magnetic flux $\Psi_{\rm open}$ is plotted in units of $\Psi_{\rm dipole\ LC}$ with the position of the Y-point. Our present result that $\Psi_{\rm open}\rightarrow 0$ as $x_{\rm Y}\rightarrow 1$ is in tension with all previous solutions of the pulsar magnetosphere. The dotted blue line corresponds to the fit $1.1\ \Psi_{\rm open}/\Psi_{\rm dipole\ LC}/x_{\rm Y}$, while the grey band corresponds to the range of values obtained from the simulations of Timokhin (2006). For example, solution No 4 from Table 1 contains the same amount of open magnetic flux as the one obtained by Hu and Beloborodov (2022), yet their Y-point lies at $x_{\rm Y}=0.85$, while ours lies at $x_{\rm Y}=0.71$. We believe that this pronounced difference is mainly due to our treatment of the separatrix current sheet as a contact discontinuity. In all previous solutions (e.g. Contopoulos et al. (1999); Spitkovsky (2006); Timokhin (2006); Hu and Beloborodov (2022); Hakobyan et al. (2023)), the separatrix contains a finite amount of poloidal magnetic flux and has a pronounced nonzero thickness. This affects the pressure balance across it which significantly modifies the position of the magnetospheric Y-point. The difference may also be partially due to our choice of a rather large stellar radius $r_{*}=R_{\rm LC}/4$ compared to previous solutions. Finally, in figure 6 the variation of the pulsar spindown energy loss $\dot{E}$ is plotted in units of $\dot{E}_{\rm vacuum}(90^{\circ})$ with the position of the Y-point. Our present result that $\dot{E}\rightarrow 0$ as $x_{\rm Y}\rightarrow 1$ is
@@
Figure 4: The distribution of $B_{z}(r,\theta=\pi/2)$ inside the Y-point for the various solutions listed in Table 1. $B_{z}$ in arbitrary units. The dotted line represents the standard $1/r^{3}$ dipolar dependence. One clearly infers both the outward stretching of the initially dipolar field (at small $r$ distances the solid lines lies below the dotted one), and the inwards squeezing of the field right behind the Y-point that is required to satisfy pressure balance. Never before has anyone seen such large increase of $B_{z}$ inside the Y-point, even in the solutions in the literature where $x_{\rm Y}\rightarrow 1$. Our results improve figure 11 of Timokhin (2006).
@@ 4 A possible connection with Fast Radio Bursts
Time-dependent numerical simulations of the pulsar magnetosphere start with an initial configuration ${\bf B}(r,\theta,\phi;t=0)={\bf B}_{\rm dipole}(r,\theta,\phi;t=0)$ that corresponds to a magnetostatic dipole in vacuum, and ${\bf E}(r,\theta,\phi;t=0)=0$ everywhere. At time $t=0$, the programmer introduces electric fields ${\bf E}(r_{*},\theta,\phi;t)=-r\sin\theta\ \Omega/c\ \hat{\phi}\ \times{\bf B}_{\rm dipole}(r_{*},\theta,\phi;t)$ along the surface of the star, thus effectively setting the central star in rotation with angular velocity $\Omega$ along the $z$-axis. Here, $r_{*}$ is the radius of the star, and ${\bf B}_{\rm dipole}(r_{*},\theta,\phi;t)$ corresponds to a dipole field that rotates with angular velocity $\Omega$ along the $z$-axis. This is the procedure followed in the pioneering paper of Spitkovsky Spitkovsky (2006) and in all time-dependent numerical simulations since then. The abrupt introduction of a non-zero electric field along the surface of the star generates a transient blast wave of poloidal electric and azimuthal magnetic fields that sweeps through the static magnetosphere. This is analogous to the transient blast wave of electric and magnetic fields generated when an electric charge is abruptly set in motion Jackson (1975) or an electric current is introduced Gourgouliatos and Lynden-Bell (2008). The aim of such simulations is to obtain the steady-state solution, thus the programmer is not interested in the blast wave and waits for it to leave the inner magnetosphere. Indeed, after that wave sweeps through the computational grid that corresponds to the inner magnetosphere, a breathing steady-state solution is established with closed and open magnetic field lines. That initial artificial polarized blast wave has the characteristics of a Fast Radio Burst (hereafter FRB).
@@
FRBs represent one of the most intriguing enigmas in the field of astrophysics today. First discovered in 2007 Lorimer et al. (2007), FRBs are transient radio pulses that originate from distant galaxies, characterized by their extremely high energy and incredibly short duration, typically lasting just a few milliseconds. Despite their fleeting nature, these cosmic phenomena release more energy in a fraction of a second than the sun does in an entire day. The exact mechanisms that produce FRBs remain a subject of intense debate and speculation among scientists. Hypotheses range from highly magnetized neutron stars, known as magnetars, to more exotic theories involving cosmic strings or even extraterrestrial intelligence. However, despite over a decade of observation and research, no single theory has been able to fully explain all observed properties of FRBs.
@@
We propose here that FRBs may be related to the electromagnetic burst of toroidal magnetic field that follows an abrupt transition from the new solution that was obtained in this paper to one of the standard magnetospheric solutions of figure Appendix 1 and Table 1. The duration of the burst will be on the order of the light crossing time of the light cylinder, namely $R_{\rm LC}/c=P/(2\pi)$, where $P$ is the spin period of the pulsar. The total energy of the burst is estimated to be equal to the difference between the electromagnetic energy of the new solution and the electromagnetic energy inside the light cylinder of a standard magnetospheric solution. An estimate of the former is obtained as follows: according to figures 1 and 4, the dipolar magnetic flux $\Psi_{\rm dipole\ LC}$ that would have extended beyond the light cylinder, is now confined to close within an annular region of area $\sim R_{\rm LC}\delta$ where $\delta\sim 0.1R_{\rm LC}$. That region has volume $\sim R_{\rm LC}^{2}\delta\sim 0.1R_{\rm LC}^{3}$, and in it $B\sim\Psi_{\rm dipole\ LC}/(0.1R_{\rm LC}^{2})\sim 10B_{*}r_{*}^{3}/R_{\rm LC}^{3}$. Therefore,
@@
Notice that in the energy estimates of this section we use a physical neutron star radius of $r_{*}\approx 10$ km. As we will now see, this is more than enough to power an FRB. In the case of an accreting recycled millisecond pulsar with surface magnetic field on the order of $B_{*}\sim 10^{8-9}$ G,
@@
where $L_{\odot}$ is the solar luminosity. This value is below the lowest limit of the low-energy FRBs presented by Gourdji et al. (2019), hence it does not seem consistent with actual FRB observations. For a newborn neutron star with a millisecond period formed via a core collapse supernova,
@@
Actual FRB energies lie between the above extreme limits, namely between $10^{37}$ and $10^{41}$ erg (e.g. Thornton et al. (2013); Amiri et al. (2021)). Therefore, we propose that FRBs correspond to millisecond pulsars with surface magnetic fields between $10^{11}$ and $10^{12}$ G.

### 11 Apr 2024 v2

@@ 3 A Spectrum of Solutions
We obtain here magnetospheric solutions with the methodology of Paper I (see Appendix A) in which the separatrix between open and closed field lines is considered to be a mathematical contact discontinuity, namely a surface of zero thickness. The tip of the closed-line region is called the Y-point. We consider in particular Y-point positions from $x_{\rm Y}\approx r_{*}/R_{\rm LC}\ll 1$ to $x_{\rm Y}\approx 1$ very close to the light cylinder. For reasons of computational convenience, we chose a rather large star with radius $r_{*}=0.25R_{\rm LC}$ (PINN methods cannot handle well large scale differences in their domain of application). We have also tried different (smaller) values of $r_{*}$ and obtained similar results. As is described in Appendix A, we first choose the angular opening of the polar cap $\theta_{\rm pc}$ from which a certain amount of magnetic flux $\Psi_{\rm open}\equiv\Psi_{\rm S}=\Psi_{\rm max}\sin^{2}\theta_{\rm pc}$ emanates and extends to infinity. This forms the so-called open-line region. In the solutions shown in Figure 3, one particular value of $\theta_{\rm pc}$, namely $\theta_{\rm pc}=1.176(r_{*}/R_{\rm LC})^{1/2}$, was specially chosen with the following in mind: it is straightforward to calculate that, in a dipolar magnetic field configuration, the magnetic field line that crosses the light cylinder corresponds to $\Psi_{\rm dipole\ LC}=\Psi_{\rm max}r_{*}/R_{\rm LC}=B_{*}r_{*}^{3}\Omega/(2c)$ (here, $B_{*}$ is the polar value of the surface magnetic field), and $\theta_{\rm dipole\ pc}=\sin^{-1}[(r_{*}/R_{\rm LC})^{1/2}]\approx(r_{*}/R_{\rm LC})^{1/2}$. In previous high-resolution solutions of the pulsar equation (e.g., Timokhin (2006)), the magnetic field line that crosses the light cylinder corresponds to $\Psi_{\rm LC\ Timokhin}=1.23\Psi_{\rm dipole\ LC}$ and $\theta_{\rm pc\ Timokhin}\approx 1.176(r_{*}/R_{\rm LC})^{1/2}$. Therefore, this particular value of $\theta_{\rm pc}$ was chosen so that the Y-point lies very close to the light cylinder. Instead, in Paper I, it was found that $R_{\rm Y}=0.88R_{\rm LC}$, which is closer to $R_{\rm dipole}(\Psi_{\rm S})=r_{*}/\sin^{2}\theta_{\rm pc}=0.81R_{\rm LC}$. This unexpected result is due to our clear treatment of the separatrix surface as a surface of zero thickness which allows the manifestation of the T-point at its tip (see Appendix B). This is not the case in all previous MHD, FFE, and PIC simulations where the separatrix surface has an unphysically large finite thickness which at its tip leads to an exterior Y-point, and a T-point some distance inside it (see Figure 3 of Contopoulos et al. (2024) for details).
@@
We also obtained the solution for $\Psi_{\rm open}=0.866\Psi_{\rm dipole\ LC}$ that corresponds to $\theta_{\rm pc}=0.97(r_{*}/R_{\rm LC})^{1/2}$ and $R_{\rm Y}=0.94R_{\rm LC}$. The reader can check that this value of $\Psi_{\rm open}$ yields a pulsar spindown rate equal to $\dot{E}=0.75\dot{E}_{\rm vacuum}(90^{\circ})$. Here, $\dot{E}_{\rm vacuum}(\lambda)$ is the spindown rate of a vacuum dipole rotator with inclination angle $\lambda$, and $\dot{E}_{\rm vacuum}(90^{\circ})=B_{*}^{2}r_{*}^{6}\Omega^{4}/(6c^{3})\equiv 2\Psi_{\rm dipole\ LC}^{2}\Omega^{2}/(3c)$. One may tentatively generalize our result for non-zero pulsar inclination angles according to Spitkovsky (2006) as $\dot{E}(\lambda)\approx 0.75\dot{E}_{\rm vacuum}(90^{\circ})(1+\sin^{2}\lambda)$, and since $\dot{E}_{\rm vacuum}(\lambda)=\dot{E}_{\rm vacuum}(90^{\circ})\sin^{2}\lambda$, we obtain that
@@
It is interesting that in all previous solutions of the FFE pulsar magnetosphere, the above ratio was found to be greater than 3 (e.g., Li et al. (2012)). This value is significantly larger than the ratio of spindown rates $\dot{E}_{\rm ON}/\dot{E}_{\rm OFF}$ observed in the intermittent pulsars PSR B1931+24, PSR J1832+0029 and PSR J1841-0500 for their corresponding ‘ON’ and ‘OFF’ states (1.5, 1.7 and 2.5, respectively, e.g., Rea et al. (2008); Wang et al. (2020)). The inability to account for observed values lower than 3 was the reason that led to the development of resistive magnetospheric solutions (e.g., Kalapotharakos et al. (2012); Li et al. (2012)). With our new solutions, it seems that there is no need for magnetospheric resistivity anymore. This result certainly merits further investigation.
Figure 3: Intermediate solutions for $\theta_{\rm pc}=\pi/2$ and $\theta_{\rm pc}=2/1.176/0.9/0.7/0\times(r_{*}/R_{\rm LC})^{1/2}$ (top left/top right/middle left/middle right/bottom left/bottom right, respectively), or equivalently $\Psi_{\rm open}=\Psi_{\rm max}$ and $\Psi_{\rm open}\equiv\Psi_{\rm S}=\Psi_{\rm max}\sin^{2}\theta_{\rm pc}=2.256/1.228/0.756/0.468/0\times\Psi_{\rm dipole\ LC}$, respectively. $\Psi$ contour levels: integer multiples of $0.1\Psi_{\rm dipole\ LC}$ with $\Psi=0$ along the $z$-axis. Red dotted lines: initial dipolar shape of the separatrix between open and closed field lines. Thin vertical line: line cylinder. Thick lines: current sheets along the separatrix and equator (and along the light cylinder in the bottom right solution).
We were able to apply our method to even smaller values of $\Psi_{\rm open}$ and $\theta_{\rm pc}$. These solutions are very hard to obtain with the methodology of Paper I because in these, the Y-point approaches very close to the light cylinder and the convergence of the solution in the open-line region between the Y-point and the light cylinder is very difficult. In the limit $\Psi_{\rm open}=\theta_{\rm pc}=0$, one reaches the solution shown in Figure 1, which was obtained independently as a solution of the pulsar equation (eq. 2) with standard methods (see Gourgouliatos and Lynden-Bell (2008)). This spectrum of solutions with smaller and smaller spindown rates is new and has never been obtained before in previous MHD, FFE, nor PIC simulations. Our results are summarized in Table 1.
@@
In Figure 4, the distribution of ${\rm B}_{z}$ is plotted along the equator inside the light cylinder and shows its clear increase right inside the Y-point at its tip, as is needed in order to satisfy Equation (13). Although this effect was known, it has never been observed so clearly before, and Figure 4 improves Figure 11 of Timokhin (2006). It is now seen clearly that, with the improved methodology of Dimitropoulos et al. (2024), the minimum ${\rm B}_{z}$ value of this sequence of solutions lies at around $x_{\rm Y}\sim 0.8$, which is near the position of the Y-point in the high-resolution PIC simulations of Cerutti et al. (2015); Hu and Beloborodov (2022); Hakobyan et al. (2023); Bransgrove et al. (2023). It may be that the Y-point naturally relaxes to that position (not at the light cylinder), as this corresponds to a minimum energy solution. This result may answer an important question in the recent literature on the pulsar magnetosphere. Note, however, that the numerical experiments in Kalapotharakos et al. (2023) lead to a different interpretation, namely that the Y-point can be pushed arbitrarily close to the stellar surface depending on the number of PIC particles that are either injected or reach the dissipative region beyond the Y-point. This is an interesting interpretation since it suggests that $x_{\rm Y}$ may decrease as the magnetosphere evolves with pulsar spindown. This will lead to a value of the braking index smaller than its canonical value of 3. Otherwise, if $x_{\rm Y}$ remains unchanged, all solutions obtained in Table 1 will have braking indices equal to 3.
@@
In Figure 5, the variation in the open magnetic flux $\Psi_{\rm open}$ is plotted in units of $\Psi_{\rm dipole\ LC}$ with the position of the Y-point. Our present result that $\Psi_{\rm open}\rightarrow 0$ as $x_{\rm Y}\rightarrow 1$ is in tension with all previous solutions of the pulsar magnetosphere. The dotted blue line corresponds to the fit $1.1\ \Psi_{\rm open}/\Psi_{\rm dipole\ LC}/x_{\rm Y}$, while the grey band corresponds to the range of values obtained from the simulations of Timokhin (2006). For example, solution No 4 from Table 1 contains the same amount of open magnetic flux as the one obtained by Hu and Beloborodov (2022), yet their Y-point lies at $x_{\rm Y}=0.85$, while ours lies at $x_{\rm Y}=0.71$. We believe that this pronounced difference is mainly due to our treatment of the separatrix current sheet as a contact discontinuity. In all previous solutions (e.g., Contopoulos et al. (1999); Spitkovsky (2006); Timokhin (2006); Hu and Beloborodov (2022); Hakobyan et al. (2023)), the separatrix contains a finite amount of poloidal magnetic flux and has a pronounced nonzero thickness. This affects the pressure balance across it which significantly modifies the position of the magnetospheric Y-point. The difference may also be partially due to our choice of a rather large stellar radius $r_{*}=R_{\rm LC}/4$ compared to previous solutions. Finally, in Figure 6 the variation in the pulsar spindown energy loss $\dot{E}$ is plotted in units of $\dot{E}_{\rm vacuum}(90^{\circ})$ with the position of the Y-point. Our present result that $\dot{E}\rightarrow 0$ as $x_{\rm Y}\rightarrow 1$ is also in tension with all previous solutions of the pulsar magnetosphere. The grey band corresponds to the spindown energy losses obtained from the simulations of Contopoulos et al. (1999); Contopoulos (2005); Timokhin (2006). As is acknowledged above, the difference between the dotted line and the grey band may be partially due to our choice of a rather large stellar radius. The value shown with the horizontal line corresponds to the particular value $\dot{E}=0.75\ \dot{E}_{\rm vacuum}(90^{\circ})$ mentioned in Equation (6).
@@
Figure 4: The distribution of ${\rm B}_{z}(r,\theta=\pi/2)$ inside the Y-point for the various solutions listed in Table 1. ${\rm B}_{z}$ is in arbitrary units. The dotted line represents the standard $1/r^{3}$ dipolar dependence. One clearly infers both the outward stretching of the initially dipolar field (at small $r$ distances the solid lines lie below the dotted one), and the inwards squeezing of the field right behind the Y-point that is required to satisfy pressure balance. Never before has anyone seen such a large increase in ${\rm B}_{z}$ inside the Y-point, even in the solutions in the literature where $x_{\rm Y}\rightarrow 1$. Our results improve Figure 11 of Timokhin (2006).
@@ 4 A Possible Connection with Fast Radio Bursts
Time-dependent numerical simulations of the pulsar magnetosphere start with an initial configuration ${\bf B}(r,\theta,\phi;t=0)={\bf B}_{\rm dipole}(r,\theta,\phi;t=0)$ that corresponds to a magnetostatic dipole in vacuum, and ${\bf E}(r,\theta,\phi;t=0)=0$ everywhere. At time $t=0$, the programmer introduces electric fields ${\bf E}(r_{*},\theta,\phi;t)=-r\sin\theta\ \Omega/c\ \hat{\phi}\ \times{\bf B}_{\rm dipole}(r_{*},\theta,\phi;t)$ along the surface of the star, thus effectively setting the central star in rotation with angular velocity $\Omega$ along the $z$-axis. Here, $r_{*}$ is the radius of the star, and ${\bf B}_{\rm dipole}(r_{*},\theta,\phi;t)$ corresponds to a dipole field that rotates with angular velocity $\Omega$ along the $z$-axis. This is the procedure followed in the pioneering paper of Spitkovsky Spitkovsky (2006) and in all time-dependent numerical simulations since then. The abrupt introduction of a non-zero electric field along the surface of the star generates a transient blast wave of poloidal electric and azimuthal magnetic fields that sweeps through the static magnetosphere. This is analogous to the transient blast wave of electric and magnetic fields generated when an electric charge is abruptly set in motion Jackson (1975) or an electric current is introduced Gourgouliatos and Lynden-Bell (2008). The aim of such simulations is to obtain the steady-state solution; thus, the programmer is not interested in the blast wave and waits for it to leave the inner magnetosphere. Indeed, after that wave sweeps through the computational grid that corresponds to the inner magnetosphere, a breathing steady-state solution is established with closed and open magnetic field lines. That initial artificial polarized blast wave has the characteristics of a fast radio burst (hereafter FRB).
@@
FRBs represent one of the most intriguing enigmas in the field of astrophysics today. First discovered in 2007 Lorimer et al. (2007), FRBs are transient radio pulses that originate from distant galaxies, characterized by their extremely high energy and incredibly short duration, typically lasting just a few milliseconds. Despite their fleeting nature, these cosmic phenomena release more energy in a fraction of a second than the sun in an entire day. The exact mechanisms that produce FRBs remain a subject of intense debate and speculation among scientists. Hypotheses range from highly magnetized neutron stars, known as magnetars, to more exotic theories involving cosmic strings or even extraterrestrial intelligence. However, despite over a decade of observation and research, no single theory has been able to fully explain all observed properties of FRBs.
@@
We propose here that FRBs may be related to the electromagnetic burst of the toroidal magnetic field that follows an abrupt transition from the new solution that was obtained in this paper to one of the standard magnetospheric solutions of Figure 1 and Table 1. The duration of the burst will be on the order of the light crossing time of the light cylinder, namely $R_{\rm LC}/c=P/(2\pi)$, where $P$ is the spin period of the pulsar. The total energy of the burst is estimated to be equal to the difference between the electromagnetic energy of the new solution and the electromagnetic energy inside the light cylinder of a standard magnetospheric solution. An estimate of the former is obtained as follows: according to Figures 1 and 4, the dipolar magnetic flux $\Psi_{\rm dipole\ LC}$ that would have extended beyond the light cylinder, is now confined to close within an annular region of area $\sim R_{\rm LC}\delta$ where $\delta\sim 0.1R_{\rm LC}$. That region has volume $\sim R_{\rm LC}^{2}\delta\sim 0.1R_{\rm LC}^{3}$, and in it, $B\sim\Psi_{\rm dipole\ LC}/(0.1R_{\rm LC}^{2})\sim 10B_{*}r_{*}^{3}/R_{\rm LC}^{3}$. Therefore,
@@
Notice that in the energy estimates of this section, we use a physical neutron star radius of $r_{*}\approx 10$ km.
As we will now see, this is more than enough to power an FRB. In the case of an accreting recycled millisecond pulsar with a surface magnetic field on the order of $B_{*}\sim 10^{8-9}$ G,
@@
where $L_{\odot}$ is the solar luminosity. This value is below the lowest limit of the low-energy FRBs presented by Gourdji et al. (2019); hence, it does not seem consistent with actual FRB observations. For a newborn neutron star with a millisecond period formed via a core collapse supernova,
@@
Actual FRB energies lie between the above extreme limits, namely between $10^{37}$ and $10^{41}$ erg (e.g., Thornton et al. (2013); Amiri et al. (2021)). Therefore, we propose that FRBs correspond to millisecond pulsars with surface magnetic fields between $10^{11}$ and $10^{12}$ G.